# Cold Jupiters and small planets: friends, foes, or indifferent?

### v1

@@ Appendix D Completenesses of the considered samples
Appendix E Cold Jupiters and their occurrence rates in the small planet and radial velocity comparison samples
Table 4: Cold Jupiters in the small planet sample.
Target $M_{\rm p}\sin(i)$ or $M_{\rm p}$ $a$ M [Fe/H] Ref.
[$\mathrm{M}_{\rm Jup}$] [AU] [M] [dex]
BD-11 4672 b $0.65_{-0.06}^{+0.05}$ $2.360\pm 0.040$ $0.650\pm 0.030$ $-0.35\pm 0.20$ Barbato et al. (2020)
GJ 676A b $5.790_{-0.477}^{+0.500}$ $1.735_{-0.060}^{+0.056}$ $0.73\pm 0.20$ $0.080$ Feng et al. (2022)
GJ 676A c $13.500_{-1.127}^{+1.000}$ $9.726_{-0.793}^{+0.629}$ $0.73\pm 0.20$ $0.080$ Feng et al. (2022)
HD 125612 b $2.974_{-0.258}^{+0.245}$ $1.349_{-0.059}^{+0.054}$ $1.11\pm 0.20$ $0.250\pm 0.040$ Feng et al. (2022)
HD 125612 d $7.180_{-0.445}^{+0.900}$ $3.982_{-0.174}^{+0.159}$ $1.11\pm 0.20$ $0.250\pm 0.040$ Feng et al. (2022)
HD 156098 c $5.228_{-1.287}^{+1.939}$ $8.41_{-0.629}^{+0.396}$ $1.29\pm 0.20$ $0.190\pm 0.100$ Feng et al. (2022)
HD 160691 b $1.67\pm 0.11$ $1.50\pm 0.02$ $1.130\pm 0.020$ $0.280\pm 0.030$ McCarthy et al. (2004)
HD 160691 c $3.10\pm 0.71$ $4.17\pm 0.07$ $1.130\pm 0.020$ $0.280\pm 0.030$ McCarthy et al. (2004)
HD 164922 b $0.344\pm 0.013$ $2.149\pm 0.025$ $0.920\pm 0.030$ $0.200\pm 0.060$ Rosenthal et al. (2021)
HD 181433 c $0.674\pm 0.003$ $1.819\pm 0.001$ $0.780$ $0.33\pm 0.10$ Horner et al. (2019)
HD 181433 d $0.612\pm 0.004$ $6.60\pm 0.22$ $0.780$ $0.33\pm 0.10$ Horner et al. (2019)
HD 190360 b $1.492\pm 0.043$ $3.955_{-0.053}^{+0.051}$ $0.990\pm 0.040$ $0.230\pm 0.060$ Rosenthal et al. (2021)
HD 204313 b $4.96_{-0.45}^{+1.70}$ $3.208_{-0.035}^{+0.034}$ $1.045\pm 0.030$ $0.180\pm 0.020$ Xiao et al. (2023)
HD 204313 e $15.300_{-5.183}^{+4.900}$ $7.457_{-0.427}^{+0.399}$ $1.045\pm 0.030$ $0.180\pm 0.020$ Feng et al. (2022)
HD 215497 c $0.35\pm 0.02$ $1.310\pm 0.020$ $0.920\pm 0.050$ $0.230\pm 0.070$ Philipot et al. (2023)
HD 219134 h $0.308\pm 0.014$ $2.968\pm 0.037$ $0.790\pm 0.030$ $0.080\pm 0.060$ Rosenthal et al. (2021)
HD 219828 c $16.10_{-1.27}^{+3.50}$ $5.657_{-0.242}^{+0.228}$ $1.18\pm 0.20$ $0.180$ Feng et al. (2022)
HD 34445 b $0.658_{-0.040}^{+0.041}$ $2.105_{-0.039}^{+0.038}$ $1.110\pm 0.060$ $0.140\pm 0.060$ Rosenthal et al. (2021)
HD 34445 g $0.38\pm 0.13$ $6.36\pm 1.02$ $1.110\pm 0.060$ $0.140\pm 0.060$ Vogt et al. (2017)
HD 39091 b/$\pi$ Men b $12.300_{-1.384}^{+1.200}$ $3.311_{-0.148}^{+0.134}$ $1.070\pm 0.040$ $0.090\pm 0.040$ Feng et al. (2022)
HD 74698 c $0.40\pm 0.06$ $4.5\pm 0.2$ $1.040\pm 0.025$ $0.07\pm 0.02$ Frensch et al. (2023)
HD 86226 b $0.45^{+0.04}_{-0.05}$ $2.730\pm 0.060$ $1.020\pm 0.060$ $0.020\pm 0.050$ Teske et al. (2020)
Kepler-48 e $2.163\pm 0.062$ $1.898$ $0.880\pm 0.060$ $0.170\pm 0.070$ Weiss et al. (2024)
Kepler-48 f $0.939\pm 0.292$ $5.719$ $0.880\pm 0.060$ $0.170\pm 0.070$ Weiss et al. (2024)
Kepler-68 d $0.749\pm 0.017$ $1.469\pm 0.010$ $1.080\pm 0.050$ $0.120\pm 0.070$ Bonomo et al. (2023)
Kepler-94 c $8.885\pm 0.117$ $1.597$ $0.810\pm 0.060$ $0.340\pm 0.070$ Weiss et al. (2024)
Kepler-129 d $8.30_{-0.70}^{+1.10}$ $4.00\pm 0.10$ $1.178\pm 0.030$ $0.290\pm 0.100$ Zhang et al. (2021)
Kepler-139 e $1.341\pm 0.134$ $3.236$ $1.078\pm 0.030$ $0.350\pm 0.040$ Weiss et al. (2024)
Kepler-407 c $11.089\pm 0.049$ $3.27$ $1.000\pm 0.060$ $0.330\pm 0.070$ Weiss et al. (2024)
Kepler-454 c $4.51\pm 0.12$ $1.287\pm 0.017$ $1.030\pm 0.040$ $0.320\pm 0.080$ Bonomo et al. (2023)
Kepler-454 d $2.31_{-0.16}^{+0.27}$ $5.10_{-0.19}^{+0.34}$ $1.030\pm 0.040$ $0.320\pm 0.080$ Bonomo et al. (2023)
K2-312 c/HD 80653 c $5.41_{-0.44}^{+0.52}$ $1.961\pm 0.027$ $1.180\pm 0.040$ $0.255\pm 0.070$ Bonomo et al. (2023)
K2-364 c/HD 137496 c $7.66\pm 0.11$ $1.2163\pm 0.0087$ $1.040\pm 0.020$ $-0.030\pm 0.040$ Azevedo Silva et al. (2022)
TOI-1669 c $0.601_{-0.076}^{+0.079}$ $1.27\pm 0.03$ $1.00\pm 0.05$ $0.26\pm 0.06$ Polanski et al. (2024)
TOI-1736 c $8.70_{-0.60}^{+1.50}$ $1.381\pm 0.017$ $1.080\pm 0.040$ $0.138\pm 0.010$ Martioli et al. (2023)
TOI-1339 f/HD 191939 f $6.50\pm 4.515$ $4.80\pm 2.20$ $0.810\pm 0.040$ $-0.150\pm 0.060$ Lubin et al. (2022)
 [^16] From left to right, the columns report the name of the cold Jupiter, its mass and semi-major axis, the mass and metallicity of the host star, and the reference for the cold-Jupiter mass determination. Asterisks denote true masses instead of minimum masses thanks to the combination of radial velocities and measurements of Hipparcos-Gaia proper motion anomalies (e.g., Kervella et al. 2022).
Table 5: Cold Jupiters in the mixed system sample.
Target $M_{\rm p}\sin(i)$ or $M_{\rm p}$ $a$ M [Fe/H] Ref.
[$\mathrm{M}_{\rm Jup}$] [AU] [M] [dex]
55 Cnc d $3.12\pm 0.10$ $5.96\pm 0.07$ $0.975\pm 0.045$ $0.383\pm 0.055$ Bourrier et al. (2018)
HIP 57274 d $0.53\pm 0.03$ $1.01$ $0.73\pm 0.05$ $0.09\pm 0.05$ Fischer et al. (2012)
Kepler-56 d $5.61\pm 0.38$ $2.16\pm 0.08$ $1.67\pm 0.15$ $0.40\pm 0.04$ Otor et al. (2016)
Kepler-88 d/KOI-142 d $3.15\pm 0.17$ $2.47$ $0.985\pm 0.021$ $0.27\pm 0.04$ Weiss et al. (2024)
TOI-4010 e $2.18_{-0.20}^{+0.21}$ $1.57_{-0.13}^{+0.12}$ $0.89\pm 0.03$ $0.37\pm 0.07$ Kunimoto et al. (2023)
WASP-47 c $1.41_{-0.07}^{+0.06}$ $1.393\pm 0.014$ $1.04\pm 0.03$ $0.38\pm 0.05$ Nascimbeni et al. (2023)
 [^17] Same as Table LABEL:table_CJs_SPsample but for cold Jupiters within mixed systems.
@@ Appendix E Cold Jupiters and their occurrence rates in the small planet and radial velocity comparison samples
Figure 8: Similarly to Fig. 1, occurrence rates of cold Jupiters with $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ in the AAT (orange), CLS (violet), and HARPS (green) radial velocity surveys, at sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. Almost identical occurrence rates are obtained for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. See also Table LABEL:table_RVsurveys.

### v2

@@ Appendix B Completenesses of the considered samples
Appendix C Cold Jupiters and their occurrence rates in the small planet and radial velocity comparison samples
Figure 7: The architectures of the systems with both SPs and CJs considered in this work and ordered according to increasing semimajor axis of the innermost planet (from bottom to top). The CJs in the left (27 systems) and right (23 systems) panels have masses $0.3\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$ and $0.5\leq{\rm M_{p}}\leq 20\,\hbox{$\mathrm{M}_{\rm Jup}$}$, respectively, with a large overlap for CJs with $0.5\leq{\rm M_{p}}\leq 13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. The planet circles have sizes proportional to their mass (see legend in the top right corner), and colors indicating orbital eccentricity (more eccentric planets are shown in yellow). Host stars are color coded as a function of stellar metallicity (metallicity increases from blue to red).
@@ Appendix C Cold Jupiters and their occurrence rates in the small planet and radial velocity comparison samples
Figure 8: Similarly to Fig. 1, occurrence rates of CJs with $M_{\rm p}=0.5-20\,\hbox{$\mathrm{M}_{\rm Jup}$}$ in the AAT (orange), CLS (violet), and HARPS (green) RV surveys, at sub-solar ($\rm[Fe/H]<-0.1$), solar ($\rm-0.1\leq[Fe/H]\leq 0.1$), and super-solar ($\rm[Fe/H]>0.1$) metallicity. Almost identical occurrence rates are obtained for $M_{\rm p}=0.3-13\,\hbox{$\mathrm{M}_{\rm Jup}$}$. See Table 17.